# TEG-DB: A Comprehensive Dataset and Benchmark of Textual-Edge Graphs

### v1

@@ 4.3 Adapting Existing Methods to Solve Problems in TEGs
PLM-based Paradigm. PLMs are trained on massive amounts of text data, allowing them to learn the semantic relationships between words, phrases, and sentences. This enables them to understand the meaning behind the text, not just on a superficial level, but also in terms of context and intent. So PLM-based methods leverage the power of PLM to enhance the text modeling within each node. The formulation of these methods is as follows:
@@
where $\boldsymbol{h}_{u}^{(k)}$ denotes the node representation of node $u$ in layer $k$ and the initial feature vector $\boldsymbol{h}_{u}^{(0)}$ of the node $u$ is obtained by encoding the text on node $u$ using the PLM. $\psi$ denotes the learnable parameters within the $\mathrm{MLP}$.
@@
where $\boldsymbol{h}_{u}^{(k)}$ denotes the node representation of node $u$ in layer $k$ and the initial node feature vector $\boldsymbol{h}_{u}^{(0)}$ uses pre-learned by PLMs or other shallow text encoder (e.g., Skip-Gram). $\boldsymbol{e}_{v,u}$ represents edge features from node $v$ to node $u$. $k$ represents the layers of GNNs, $\mathcal{N}$ denotes the set of neighbors, $u$ denotes the target node, $\boldsymbol{\omega}$ means the learning parameters in GNNs.
However, this approach presents two primary issues: (1) In TEGs, the text on edges is highly informative, yet most GNN methods neglect edge features. This omission can limit the model’s capacity to capture more complex relationships within the graph, thereby reducing its effectiveness in accurately representing the intricate interdependencies and interactions between nodes. (2) GNN-based methods are unable to fully capture the contextualized text semantics of edges. This is because the textual information associated with edges can be highly variable and context-dependent, making it challenging to represent and incorporate effectively within the GNN framework.
@@ 5 Experiments
Methods Goodreads-Children Goodreads-Crime GPT-3.5-TURBO BERT-Large BERT-Base w/o edge text GPT-3.5-TURBO BERT-Large BERT-Base w/o edge text AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 MLP 0.8952 0.8198 0.8948 0.8193 0.8947 0.8192 0.8929 0.8181 0.8911 0.8144 0.8909 0.8145 0.8920 0.8153 0.8913 0.8149 GraphSAGE 0.9520 0.8866 0.9493 0.8821 0.9503 0.8848 0.9400 0.8736 0.9241 0.8541 0.9537 0.8887 0.9529 0.8868 0.9053 0.8320 GeneralConv 0.9519 0.8907 0.9521 0.8921 0.9540 0.8953 0.9356 0.8735 0.9325 0.8625 0.9568 0.8957 0.9257 0.8526 0.9117 0.8426 GINE 0.9518 0.8939 0.9463 0.8878 0.9491 0.8914 0.9389 0.8748 0.9125 0.8429 0.9517 0.8878 0.9538 0.8928 0.9132 0.8448 EdgeConv 0.9487 0.8851 0.9488 0.8884 0.9504 0.8891 0.9352 0.8765 0.9104 0.8410 0.9545 0.8914 0.9535 0.8897 0.9036 0.8345 GraphTransformer 0.9487 0.8751 0.9441 0.8742 0.9431 0.8763 0.9241 0.8333 0.9078 0.8309 0.9465 0.8769 0.9479 0.8817 0.8985 0.8256 Methods Amazon-Apps Amazon-Movie GPT-3.5-TURBO BERT-Large BERT-Base w/o edge text GPT-3.5-TURBO BERT-Large BERT-Base w/o edge text AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 MLP 0.8642 0.7752 0.8639 0.7698 0.8634 0.7698 0.8655 0.7738 0.8227 0.7269 0.8213 0.7553 0.8349 0.7555 0.8205 0.7317 GraphSAGE 0.8662 0.7853 0.8813 0.7971 0.8783 0.8015 0.8634 0.7366 0.8500 0.7665 0.9067 0.8298 0.9178 0.8426 0.8507 0.7591 GeneralConv 0.8810 0.8178 0.8768 0.8131 0.8757 0.8090 0.8680 0.8129 0.8659 0.7928 0.9206 0.8485 0.8937 0.8483 0.8617 0.7918 GINE 0.8559 0.8099 0.8680 0.8092 0.8555 0.8123 0.8671 0.8065 0.8603 0.7911 0.9187 0.8454 0.9165 0.8456 0.8591 0.7879 EdgeConv 0.8720 0.8180 0.8813 0.8153 0.8804 0.8184 0.8520 0.8043 0.8565 0.7842 0.9171 0.8436 0.9181 0.8468 0.8552 0.7837 GraphTransformer 0.8395 0.7647 0.8748 0.7926 0.8736 0.7846 0.8469 0.7329 0.8339 0.7453 0.9035 0.8196 0.9044 0.8185 0.8393 0.7550 Methods Citation Twitter GPT-3.5-TURBO BERT-Large BERT-Base w/o edge text GPT-3.5-TURBO BERT-Large BERT-Base w/o edge text AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 ACC F1 AUC F1 MLP 0.9170 0.8598 0.9173 0.8561 0.8935 0.8613 0.8857 0.8015 0.6991 0.5430 0.8115 0.7898 0.8136 0.7148 0.7007 0.5430 GraphSAGE 0.9369 0.8758 0.9457 0.8832 0.9780 0.9300 0.8925 0.8345 0.6779 0.6193 0.8609 0.8177 0.8359 0.7964 0.5668 0.5940 GeneralConv 0.9258 0.8739 0.9281 0.8637 0.9327 0.8757 0.8984 0.8397 0.7888 0.7094 0.8531 0.7756 0.8062 0.6552 0.7017 0.6163 GINE 0.9482 0.8939 0.9443 0.8825 0.9736 0.9272 0.8744 0.8145 0.6696 0.6135 0.8306 0.7719 0.8738 0.7880 0.7213 0.6161 EdgeConv 0.7136 0.5393 0.7132 0.5352 0.7401 0.6526 0.6965 0.5449 0.6854 0.6123 0.8290 0.6614 0.7513 0.6745 0.6124 0.5664 GraphTransformer 0.9350 0.8697 0.9439 0.8713 0.9789 0.9320 0.9172 0.8441 0.6859 0.6764 0.8967 0.8223 0.8768 0.8165 0.5908 0.5423
@@ 5.1 Experimental Settings
Baselines. (1) For the PLM-based Paradigm, we use three various sizes of PLM to encode texts in nodes for generating initial embeddings for nodes. These three models, representing large, medium, and small scales, include GPT-3.5-TURBO as the large-scale model, Bert-Large [8] as the medium-scale model, and Bert-Base [8] as the small-scale model. (2) For GNN-based methods, we select 5 popular GNN models: GraphSAGE [11], GeneralConv [44], GINE [14], EdgeConv [39], and GraphTransformer [45]. We utilize three distinct scales of the PLMs, which are identical to those employed in the PLM-based paradigm, to encode text in nodes and edges. Afterward, these text embeddings on nodes and edges serve as their initial characteristics. (3) For LLM-based Predictor methods, we chose state-of-the-art models GPT-3.5-TURBO and GPT-4, accessed via API, balancing performance and cost considerations.
Implementation details. We conduct experiments on 3 PLM-based, 18 GNN-based, and 2 LLM-based methods. For PLM-based methods, the dimensions of node embedding are 3072, 1024, and 768 generated by GPT-3.5-TURBO, Bert-Large, and Bert respectively. We set the MLP hidden layer to 2, with the number of hidden units in each layer being one-fourth of the units in the previous layer. For GNN-based methods, we adhere to the settings outlined in the respective paper. The parameters shared by all GNN models include dimensions of node and edge embeddings, model layers, and hidden units, with respective values set to 3072, 1024, and 768, as generated by GPT-3.5-TURBO, Bert-Large, and Bert, and 2, 256, respectively. We utilize cross-entropy loss with the Adam optimizer to train and optimize all the above models. The batch size is 1024. Each experiment is repeated three times. GNNs are mainly derived from the implementation in the PyG library [9]. For the node classification task, numerical node labels corresponding to the nodes within the graph are necessary. This involves converting the categorical node categories found in the original data into numerical node labels within the graph. For the link prediction, we randomly sample node pairs that do not exist in the graph as negative samples, along with some edges present as positive samples. For LLM-based predictor methods, we focus on node classification and link prediction tasks. For node classification, inspired by the recent LLM-based classification algorithm [26], we use GPT-4 and GPT-3.5-TURBO models to predict the classification of text nodes by providing the probability for each class. We randomly select 1,000 text nodes along with all classification labels for this task. For the link prediction task, we also apply the GPT-4 and GPT-3.5-TURBO models to determine whether two text edges are related, providing an answer with the corresponding probability. For this task, we randomly select 1,000 pairs of positive text edge indices from the graph and an equal number of negative edges.
@@
Methods Goodreads-Children Goodreads-Crime GPT-3.5-TURBO BERT-Large BERT w/o edge text GPT-3.5-TURBO BERT-Large BERT w/o edge text AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* MLP 0.8505 0.5663 0.8593 0.5810 0.8597 0.5749 0.8452 0.5811 0.9149 0.6615 0.9150 0.6619 0.9151 0.6602 0.9154 0.6624 GraphSAGE 0.9342 0.7871 0.9162 0.7497 0.9152 0.7440 0.8713 0.6227 0.9549 0.8189 0.9445 0.7832 0.9463 0.7848 0.9221 0.7048 GeneralConv 0.9352 0.7846 0.9161 0.7502 0.9152 0.7451 0.8681 0.6162 0.9546 0.8200 0.9446 0.7854 0.9456 0.7888 0.9225 0.7262 GINE 0.9324 0.7777 0.9154 0.7466 0.9137 0.7552 0.8523 0.6558 0.9504 0.8073 0.9410 0.7766 0.9429 0.7852 0.9155 0.7117 EdgeConv 0.9338 0.7808 0.9128 0.7463 0.9121 0.7452 0.8583 0.6466 0.9490 0.8052 0.9400 0.7657 0.9405 0.7726 0.9187 0.6830 GraphTransformer 0.9340 0.7823 0.9137 0.7497 0.9150 0.7491 0.8517 0.6565 0.9505 0.8151 0.9452 0.7795 0.9464 0.7834 0.9220 0.6944 Methods Amazon-Apps Amazon-Movie GPT-3.5-TURBO BERT-Large BERT w/o edge text GPT-3.5-TURBO BERT-Large BERT w/o edge text AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* MLP 0.7520 0.3204 0.8935 0.4169 0.8970 0.3107 0.7352 0.3067 0.9618 0.5279 0.9752 0.5331 0.9750 0.5173 0.9493 0.4625 GraphSAGE 0.9274 0.3899 0.9226 0.3794 0.9229 0.3929 0.9161 0.3348 0.9674 0.5165 0.9773 0.4919 0.9771 0.5185 0.9681 0.5096 GeneralConv 0.8947 0.3604 0.9171 0.3817 0.9223 0.3803 0.9151 0.3932 0.9775 0.5156 0.9768 0.4827 0.9768 0.5006 0.9757 0.5115 GINE 0.9170 0.3588 0.9170 0.2623 0.9185 0.3592 0.9028 0.3507 0.9507 0.4246 0.9758 0.4781 0.9759 0.5085 0.9168 0.4127 EdgeConv 0.8764 0.3477 0.8639 0.2739 0.8800 0.3063 0.8568 0.2247 0.9360 0.5060 0.9372 0.4672 0.9263 0.4743 0.9492 0.4853 GraphTransformer 0.9195 0.3548 0.9217 0.3425 0.9225 0.3818 0.9155 0.3860 0.9763 0.5175 0.9764 0.4856 0.9771 0.5124 0.9756 0.5126 Methods Citation Twitter GPT-3.5-TURBO BERT-Large BERT w/o edge text GPT-3.5-TURBO BERT-Large BERT w/o edge text ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 MLP 0.7868 0.7859 0.7515 0.7471 0.8044 0.8032 0.7293 0.7271 0.8115 0.7261 0.8361 0.8193 0.8533 0.8329 0.8196 0.7383 GraphSAGE 0.7883 0.7874 0.7559 0.7525 0.8046 0.8060 0.7341 0.7308 0.8411 0.7903 0.8446 0.8305 0.8384 0.8247 0.8286 0.7802 GeneralGNN 0.7906 0.7889 0.7546 0.7526 0.8057 0.8042 0.7361 0.7337 0.8610 0.8397 0.8368 0.8131 0.8609 0.8513 0.8401 0.8089 GINE 0.7934 0.7925 0.7599 0.7574 0.8106 0.8100 0.7316 0.7284 0.8438 0.8186 0.8401 0.8255 0.8460 0.8328 0.8254 0.7907 EdgeConv 0.4140 0.3845 0.4082 0.3763 0.4200 0.3906 0.3935 0.3541 0.8551 0.8442 0.8649 0.8574 0.8694 0.8607 0.8529 0.8431 GraphTransformer 0.7903 0.7885 0.7531 0.7517 0.8070 0.8056 0.7369 0.7351 0.8563 0.8273 0.8342 0.8211 0.8402 0.8261 0.8197 0.7888

### v3

@@ 4.3 Adapting Existing Methods to Solve Problems in TEGs
PLM-based Paradigm. PLMs are trained on massive amounts of text data, allowing them to learn the semantic relationships between words, phrases, and sentences. This enables them to understand the meaning behind the text, not just on a superficial level, but also in terms of context and intent. So PLM-based methods leverage the power of PLM to enhance the text modeling within each node and edge, along with an extra multilayer perception (MLP) to integrate their textual information from TEG. The formulation of these methods is as follows:
@@
$\boldsymbol{h}_{u}^{(0)}$ $=\operatorname{PLM}(T_{u})+\sum_{v\in\mathcal{N}(u)}\operatorname{PLM}(T_{e_{v,u}})$
where $\boldsymbol{h}_{u}^{(k)}$ denotes the node representation of node $u$ in layer $k$ of Multilayer Perceptron (MLP). $T_{u}$ and $T_{e_{v,u}}$ represent the raw text on node $u$ and edge $e_{v,u}$, respectively. The initial feature vector $\boldsymbol{h}_{u}^{(0)}$ of node $u$ is derived by encoding the text on node $u$ and its neighboring edges using the Pre-trained Language Model (PLM). $\mathcal{N}$ denotes the set of neighbors. $\psi$ refers to the trainable parameters within the MLP.
@@
where $\boldsymbol{h}_{u}^{(k)}$ denotes the node representation of node $u$ in layer $k$ of GNN and the initial node feature vector $\boldsymbol{h}_{u}^{(0)}$ is obtained by embedding its raw text through PLMs. $e_{v,u}$ denotes the edge from node $v$ to node $u$ and its features $\boldsymbol{e}_{v,u}$ are likewise derived from PLMs based on its raw text embeddings. $k$ represents the layers of GNNs, $\mathcal{N}$ denotes the set of neighbors, $u$ denotes the target node, $\boldsymbol{\omega}$ means the learning parameters in GNNs.
However, this approach presents two primary issues: (1) Existing Graph ML methods like GNNs typically work on structured attributes on edges instead of texts Jin et al. (2023c). In TEGs, edges are texts that contain rich semantic information, which is way beyond the typical focus of GNNs that are commonly based on connectivity (i.e., binary attribute denoting whether there is a connection or not) and edge attributes (i.e., categorical or numerical values on the edges). (2) GNN-based methods are limited in capturing the contextualized semantics of edge texts Yan et al. (2023). In TEGs, where edge and node texts are often entangled, converting them into separate node and edge embeddings during the embedding process can result in the loss of critical information about their interdependence, which diminishes the effectiveness of GNNs throughout the entire message-passing process.
Entangled GNN-based Paradigm. Traditional edge-aware GNN-based approaches that first learn edge text embeddings and then apply GNNs have limitations for TEG data because edge texts and node texts are often closely entangled. Separating them into distinct node and edge embeddings may impair important information regarding their interaction. For instance, in a citation graph where each node represents a paper, an edge might indicate that one paper cites, criticizes, or utilizes a specific part of another paper. Therefore, the edge does not represent the relationship between the entirety of the two nodes, posing a significant challenge for methods that rely on node or edge embeddings representing the entirety of a node or edge. To avoid information loss during the interaction between nodes and edges after text embedding, we propose an approach that first entangles the edge text and node text before performing the embedding. The embedding obtained in this way is then added to the message-passing operation for each pair of connected nodes. The formulation of these methods is as follows:
$\boldsymbol{h}_{u}^{(k+1)}$ $=\operatorname{UPDATE}_{\boldsymbol{\omega}}^{(k)}\left(\boldsymbol{h}_{u}^{(k)},\operatorname{AGGREGATE}_{\boldsymbol{\omega}}^{(k)}\left(\left\{\boldsymbol{h}_{v}^{(k)},v\in\mathcal{N}(u)\right\}\right)\right)$ (3)
$\boldsymbol{h}_{u}^{0}$ $=\operatorname{PLM}(T_{u},\{T_{v},T_{{e}_{v,u}},v\in\mathcal{N}(u)\})$
where $\boldsymbol{h}_{u}^{(k)}$ denotes the node representation of node $u$ in layer $k$ of the GNN. $T_{v}$, $T_{u}$, and $T_{e_{v,u}}$ represent the raw text on node $v$, node $u$, and the edge from $v$ to $u$, respectively. The initial node feature vector $\boldsymbol{h}_{u}^{(0)}$ is obtained by embedding the entangled raw text of node $u$ and its neighborhood through PLMs. $k$ represents the layers of GNNs, $\mathcal{N}$ denotes the set of neighbors, $u$ denotes the target node, $\boldsymbol{\omega}$ means the learning parameters in GNNs.
The advantage of this method over existing approaches is its ability to effectively preserve the semantic relationships between nodes and edges, making it more suitable for capturing complex relationships.
@@
Methods Children Crime Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 MLP 0.9146 0.8459 0.8952 0.8198 0.8948 0.8193 0.8947 0.8192 0.8929 0.8181 0.9030 0.8429 0.8911 0.8144 0.8909 0.8145 0.8920 0.8153 0.8913 0.8149 GraphSAGE 0.9744 0.9011 0.9520 0.8866 0.9493 0.8821 0.9503 0.8848 0.9400 0.8736 0.9331 0.8629 0.9241 0.8541 0.9537 0.8887 0.9529 0.8868 0.9053 0.8320 General GNN 0.9653 0.9015 0.9519 0.8907 0.9521 0.8921 0.9540 0.8953 0.9356 0.8735 0.9356 0.8792 0.9325 0.8625 0.9568 0.8957 0.9257 0.8526 0.9117 0.8426 GINE 0.9558 0.9132 0.9518 0.8939 0.9463 0.8878 0.9491 0.8914 0.9389 0.8748 0.9324 0.8589 0.9125 0.8429 0.9517 0.8878 0.9538 0.8928 0.9132 0.8448 EdgeGNN 0.9604 0.9055 0.9487 0.8851 0.9488 0.8884 0.9504 0.8891 0.9352 0.8765 0.9309 0.8575 0.9104 0.8410 0.9545 0.8914 0.9535 0.8897 0.9036 0.8345 GraphTransformer 0.9625 0.8950 0.9487 0.8751 0.9441 0.8742 0.9431 0.8763 0.9241 0.8333 0.9123 0.8592 0.9078 0.8309 0.9465 0.8769 0.9479 0.8817 0.8985 0.8256 Methods Amazon-Apps Amazon-Movie Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 MLP 0.8950 0.7980 0.8642 0.7752 0.8639 0.7698 0.8634 0.7698 0.8655 0.7738 0.8509 0.7490 0.8227 0.7269 0.8349 0.7553 0.8349 0.7555 0.8205 0.7317 GraphSAGE 0.8911 0.8073 0.8662 0.7853 0.8813 0.7971 0.8783 0.8015 0.8634 0.7366 0.8725 0.7911 0.8500 0.7665 0.9067 0.8298 0.9178 0.8426 0.8507 0.7591 General GNN 0.8956 0.8340 0.8810 0.8178 0.8768 0.8131 0.8757 0.8090 0.8680 0.8129 0.8849 0.8134 0.8659 0.7928 0.9206 0.8485 0.8937 0.8483 0.8617 0.7918 GINE 0.8875 0.8179 0.8559 0.8099 0.8680 0.8092 0.8555 0.8123 0.8671 0.8065 0.8712 0.8154 0.8603 0.7911 0.9187 0.8454 0.9165 0.8456 0.8591 0.7879 EdgeGNN 0.8956 0.8403 0.8720 0.8180 0.8813 0.8153 0.8804 0.8184 0.8520 0.8043 0.8708 0.8035 0.8565 0.7842 0.9171 0.8436 0.9181 0.8468 0.8552 0.7837 GraphTransformer 0.8634 0.7820 0.8395 0.7647 0.8748 0.7926 0.8736 0.7846 0.8469 0.7329 0.8537 0.7698 0.8339 0.7453 0.9035 0.8196 0.9044 0.8185 0.8393 0.7550 Methods Citation Twitter Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 MLP 0.9251 0.8679 0.9170 0.8598 0.9173 0.8561 0.8935 0.8613 0.8857 0.8015 0.7085 0.5669 0.6991 0.5430 0.8115 0.7898 0.8136 0.7148 0.7007 0.5430 GraphSAGE 0.9494 0.8972 0.9369 0.8758 0.9457 0.8832 0.9780 0.9300 0.8925 0.8345 0.6998 0.6486 0.6779 0.6193 0.8609 0.8177 0.8359 0.7964 0.5668 0.5940 General GNN 0.9470 0.8840 0.9258 0.8739 0.9281 0.8637 0.9327 0.8757 0.8984 0.8397 0.8118 0.7247 0.7888 0.7094 0.8531 0.7756 0.8062 0.6552 0.7017 0.6163 GINE 0.9538 0.9085 0.9482 0.8939 0.9443 0.8825 0.9736 0.9272 0.8744 0.8145 0.6835 0.6345 0.6696 0.6135 0.8306 0.7719 0.8738 0.7880 0.7213 0.6161 EdgeGNN 0.7382 0.5545 0.7136 0.5393 0.7132 0.5352 0.7401 0.6526 0.6965 0.5449 0.6940 0.6214 0.6854 0.6123 0.8290 0.6614 0.7513 0.6745 0.6124 0.5664 GraphTransformer 0.9536 0.8963 0.9350 0.8697 0.9439 0.8713 0.9789 0.9320 0.9172 0.8441 0.7030 0.6824 0.6859 0.6764 0.8967 0.8223 0.8768 0.8165 0.5908 0.5423
@@ 5.1 Experimental Settings
Baselines. (1) For the PLM-based Paradigm, we use three various sizes of PLM to encode texts in nodes for generating initial embeddings for nodes. These three models, representing large, medium, and small scales, include GPT-3.5-TURBO as the large-scale model, Bert-Large Devlin et al. (2018) as the medium-scale model, and Bert-Base Devlin et al. (2018) as the small-scale model. (2) For Edge-aware GNN-based methods, we select 5 popular GNN models: GraphSAGE Hamilton et al. (2017), GeneralConv You et al. (2020), GINE Hu et al. (2020b), EdgeGNN Wang et al. (2019), and GraphTransformer Yun et al. (2019). We utilize three distinct scales of the PLMs, which are identical to those employed in the PLM-based paradigm, to encode text in nodes and edges. Afterward, these text embeddings on nodes and edges serve as their initial characteristics. (3) In the Entangled GNN-based approaches, the experimental setting is similar to Edge-aware GNN-based methods, with the key difference being the utilization of GPT-3.5-TURBO as the PLM to encode raw text in nodes and edges in an entangled manner. (4) For LLM-based Predictor methods, we chose state-of-the-art models GPT-3.5-TURBO and GPT-4, accessed via API, balancing performance and cost considerations.
Implementation details. We conduct experiments on 3 PLM-based, 18 GNN-based, and 2 LLM-based methods. For PLM-based methods, the dimensions of node embedding are 3072, 1024, and 768 generated by GPT-3.5-TURBO, Bert-Large, and Bert respectively. We set the MLP hidden layer to 2, with the number of hidden units in each layer being one-fourth of the units in the previous layer. For GNN-based methods, we adhere to the settings outlined in the respective paper. The parameters shared by all GNN models include dimensions of node and edge embeddings, model layers, and hidden units, with respective values set to 3072, 1024, and 768, as generated by GPT-3.5-TURBO, Bert-Large, and Bert, and 2, 256, respectively. We utilize cross-entropy loss with the Adam optimizer to train and optimize all the above models. The batch size is 1024. Each experiment is repeated three times. See Appendix B.1 for more details.
@@
Method Children Crime Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* MLP 0.8785 0.5904 0.8505 0.5663 0.8593 0.5810 0.8597 0.5749 0.8452 0.5811 0.9253 0.6842 0.9149 0.6615 0.9150 0.6619 0.9151 0.6602 0.9154 0.6624 GraphSAGE 0.9569 0.8041 0.9342 0.7871 0.9162 0.7497 0.9152 0.7440 0.8713 0.6227 0.9663 0.8325 0.9549 0.8189 0.9445 0.7832 0.9463 0.7848 0.9221 0.7048 General GNN 0.9534 0.7942 0.9352 0.7846 0.9161 0.7502 0.9152 0.7451 0.8681 0.6162 0.9732 0.8437 0.9546 0.8200 0.9446 0.7854 0.9456 0.7888 0.9225 0.7262 GINE 0.9529 0.7930 0.9324 0.7777 0.9154 0.7466 0.9137 0.7552 0.8523 0.6558 0.9636 0.8260 0.9504 0.8073 0.9410 0.7766 0.9429 0.7852 0.9155 0.7117 EdgeGNN 0.9542 0.7890 0.9338 0.7808 0.9128 0.7463 0.9121 0.7452 0.8583 0.6466 0.9581 0.8179 0.9490 0.8052 0.9400 0.7657 0.9405 0.7726 0.9187 0.6830 GraphTransformer 0.9525 0.7902 0.9340 0.7823 0.9137 0.7497 0.9150 0.7491 0.8517 0.6565 0.9613 0.8322 0.9505 0.8151 0.9452 0.7795 0.9464 0.7834 0.9220 0.6944 Method Amazon-Apps Amazon-Movie Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* AUC* F1* MLP 0.7750 0.3429 0.7520 0.3204 0.8935 0.4169 0.8970 0.3107 0.7352 0.3067 0.9736 0.5475 0.9618 0.5279 0.9752 0.5331 0.9750 0.5173 0.9493 0.4625 GraphSAGE 0.9439 0.4114 0.9274 0.3899 0.9226 0.3794 0.9229 0.3929 0.9161 0.3348 0.9764 0.5325 0.9674 0.5165 0.9773 0.4919 0.9771 0.5185 0.9681 0.5096 General GNN 0.9138 0.3806 0.8947 0.3604 0.9171 0.3817 0.9223 0.3803 0.9151 0.3932 0.9969 0.5301 0.9775 0.5156 0.9768 0.4827 0.9768 0.5006 0.9757 0.5115 GINE 0.9356 0.3862 0.9170 0.3588 0.9170 0.2623 0.9185 0.3592 0.9028 0.3507 0.9732 0.4531 0.9507 0.4246 0.9758 0.4781 0.9759 0.5085 0.9168 0.4127 EdgeGNN 0.8857 0.3749 0.8764 0.3477 0.8639 0.2739 0.8800 0.3063 0.8568 0.2247 0.9483 0.5224 0.9360 0.5060 0.9372 0.4672 0.9263 0.4743 0.9492 0.4853 GraphTransformer 0.9400 0.3772 0.9195 0.3548 0.9217 0.3425 0.9225 0.3818 0.9155 0.3860 0.9910 0.5285 0.9763 0.5175 0.9764 0.4856 0.9771 0.5124 0.9756 0.5126 Method Citation Twitter Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None Entangled-GPT GPT-3.5-TURBO BERT-Large BERT None ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 MLP 0.7892 0.7879 0.7868 0.7859 0.7515 0.7471 0.8044 0.8032 0.7493 0.7471 0.8253 0.7549 0.8115 0.7261 0.8361 0.8193 0.8533 0.8329 0.8196 0.7383 GraphSAGE 0.7984 0.8144 0.7883 0.7874 0.7559 0.7525 0.8046 0.8060 0.7341 0.7308 0.8614 0.8055 0.8411 0.7903 0.8446 0.8305 0.8384 0.8247 0.8286 0.7802 General GNN 0.8079 0.8042 0.7906 0.7889 0.7546 0.7526 0.8057 0.8042 0.7361 0.7337 0.8725 0.8574 0.8610 0.8397 0.8368 0.8131 0.8609 0.8513 0.8401 0.8089 GINE 0.8055 0.8141 0.7934 0.7925 0.7599 0.7574 0.8106 0.8100 0.7316 0.7284 0.8649 0.8386 0.8438 0.8186 0.8401 0.8255 0.8460 0.8328 0.8254 0.7907 EdgeGNN 0.4261 0.3957 0.4140 0.3845 0.4082 0.3763 0.4200 0.3906 0.3935 0.3541 0.8714 0.8530 0.8551 0.8442 0.8649 0.8574 0.8694 0.8607 0.8529 0.8431 GraphTransformer 0.8022 0.7944 0.7903 0.7885 0.7531 0.7517 0.8070 0.8056 0.7369 0.7351 0.8720 0.8369 0.8563 0.8273 0.8342 0.8211 0.8402 0.8261 0.8197 0.7888